Bexley:

The circumference of a circular patio is 52.38 feet. What is the area of the patio? Use 3.14 for an approximation of pi. Then round to the nearest tenth.

phoebe:

circumference is 2*pi*r 53.38 (which is ur circumference) = 2*3.14*r 53.38 = 6.28r r = 53.38/6.28 r = 8.5 ft i hope that helps!

phoebe:

thats the radius, which then u can use the formula of pi r^

Bexley:

@phoebe wrote:
circumference is 2*pi*r 53.38 (which is ur circumference) = 2*3.14*r 53.38 = 6.28r r = 53.38/6.28 r = 8.5 ft i hope that helps!
Thank you! I don't really understand how people can do this type of math it kinda seems hard to me